OpenStudy (anonymous):

What is the slope of a line that runs parallel to y = -x + 7? Use a number to fill in the blank.

OpenStudy (anonymous):

slope of \(y=mx+b\) is \(m\) slope of \(y=-x+7\) is ?

OpenStudy (anonymous):

I don't get it... .-.

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

that doesn't really help much .-.

OpenStudy (anonymous):

your real job here is to find the slope of \(y=-x+7\) because the slope of a parallel line is the same

OpenStudy (anonymous):

so the question is really this "what is the slope of \(y=-x+7\)?"

OpenStudy (anonymous):

Slope of lines that are parallel are the same So your slope in slope intercept y=mx+b (y=-x+7) is -1

OpenStudy (anonymous):

and you can visualize this as \[y=-1\times x+7\]

OpenStudy (anonymous):

What @satellite73 said :)

OpenStudy (anonymous):

match this with \[y=mx+b\]and you see that \(m=-1\)

OpenStudy (anonymous):

hope it is clearer now